OpenStudy (anonymous):

My textbook says M(x,y)dx+N(x,y)dy=0 is homogenous if M(tx,ty)=tαM(x,y) and N(tx,ty)=tαN(x,y) . What does that mean? I know that if the function equals F=x/y then it's homogenous. How do those two rules relate?

OpenStudy (amistre64):

homogenous .... cant recall what it means; but its got the property that it can be scaled

OpenStudy (amistre64):

x+y^2 is not homogenous because when we include the "t" we get: tx + t^2y^2 t(x+ty) is not the same

OpenStudy (anonymous):

Are you talking about differential equations?

OpenStudy (amistre64):

no; just a property of homogenoues and nonhomogenous equations; i cant recall how that relates into differential equations tho

OpenStudy (anonymous):

When you wrote tx+t^2y^2, should that be tx+ty^2?

OpenStudy (amistre64):

almost: prolly more proper like this: tx + (ty)^2

OpenStudy (amistre64):

the t helps to visuallize the homogenous property. If we can pull out the "t" and the equation remains the same; it is considered to be homogoneus

OpenStudy (amistre64):

x/y is homogen.. because: tx/ty = x/y ... so the equation remains unchanged when we test for t

OpenStudy (amistre64):

So in relation to the question: the diff equation is homogen if M and N are homogen

OpenStudy (anonymous):

My book says that x^3 +y^3 +1 isn't homogenous. Is that because if you put a "t" by the x and y, you can't take it out by saying t^3(x,y) becasue then the 1 will be a t^3, which isn't true?

OpenStudy (amistre64):

the +1 prolly messes it up; yeah

OpenStudy (anonymous):

I think it's easier (for me, at any rate) just to say that if when you put (whatever) in derivative form, there is some way of writing it as a function y/x.

OpenStudy (anonymous):

It definnitely is but my professor said we might have to show the other way on a exam :-)

OpenStudy (anonymous):

he forgot the "n" so it doesn't count lol anyways thanks

OpenStudy (anonymous):

Heh, if you delete it, it didn't happen....

OpenStudy (amistre64):

:) what didnt happen

OpenStudy (anonymous):

I guess it's the same thing really, the t has to depend on y/x anyway. f(x,y) = f(1/x x, 1/x y) = f(1,y/x) -> f(x,y) = h(y/x).